# Bilevel ZOFO: Bridging Parameter-Efficient and Zeroth-Order Techniques for Efficient LLM Fine-Tuning and Meta-Training

### v1

@@ Abstract
Fine-tuning pre-trained Large Language Models (LLMs) for downstream tasks using First-Order (FO) optimizers presents significant computational challenges. Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed to address these challenges by freezing most model parameters and training only a small subset. While PEFT is efficient, it may not outperform full fine-tuning when high task-specific performance is required. Zeroth-Order (ZO) methods offer an alternative for fine-tuning the entire pre-trained model by approximating gradients using only the forward pass, thus eliminating the computational burden of back-propagation in first-order methods. However, when implementing ZO methods, a hard prompt is crucial, and relying on simple, fixed hard prompts may not be optimal. In this paper, we propose a bilevel optimization framework that complements ZO methods with PEFT to mitigate sensitivity to hard prompts while efficiently and effectively fine-tuning LLMs. Our Bilevel ZOFO (Zeroth-Order-First-Order) method employs a double-loop optimization strategy, where only the gradient of the PEFT model and the forward pass of the base model are required. We provide convergence guarantees for Bilevel ZOFO. Empirically, we demonstrate that Bilevel ZOFO outperforms both PEFT and ZO methods in single-task settings while maintaining similar memory efficiency. Additionally, we show its strong potential for multitask learning. Compared to current first-order meta-training algorithms for multitask learning, our method has significantly lower computational demands while maintaining or improving performance.
Machine Learning, ICML
@@ 1 Introduction
Fine-tuning pretrained Large Language Models (LLMs) has become a standard approach for downstream tasks. Traditional first-order (FO) optimizers like Adam (Kingma & Ba, 2015), commonly used for this process, rely on backpropagation. However, as highlighted in Malladi et al. (2023), computing gradients for LLMs can require up to 12 times the memory needed for inference. This scaling challenge becomes even more pronounced as models grow larger, imposing significant memory demands and complicating the fine-tuning process, especially in resource-constrained environments.
@@
To address these computational challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have been developed. These techniques freeze most of the model’s parameters and train only a small subset, significantly reducing both memory and computational overhead. Popular PEFT approaches include prompt tuning, LoRA fine-tuning, and prefix tuning. Prompt tuning (Lester et al., 2021; Qin & Eisner, 2021; Yu et al., 2023) optimizes continuous prompt vectors that are concatenated with the input embeddings, while prefix tuning (Li & Liang, 2021) introduces learnable prefix tokens that serve as conditioning variables at each transformer layer. LoRA (Low-Rank Adaptation) (Hu et al., 2022; Houlsby et al., 2019) modifies the model’s attention and feedforward layers by injecting low-rank trainable matrices, further reducing the resources required for fine-tuning.
@@
While Parameter-Efficient Fine-Tuning (PEFT) methods reduce training costs and memory usage, they may not always achieve the same level of task-specific performance as full model fine-tuning. Research has shown that for tasks requiring high accuracy, complex adaptations, or domain-specific knowledge, full fine-tuning often outperforms PEFT approaches due to its ability to adjust all model parameters for better adaptation (Hu et al., 2022; Li & Liang, 2021; Zaken et al., 2022). To make full model fine-tuning more computationally feasible, zeroth-order methods offer an alternative by reducing the high computational cost. Rather than computing gradients via backpropagation, zeroth-order methods estimate the gradient using only the forward pass. Initially explored in the 1990s (Spall, 1992; Nesterov & Spokoiny, 2017; Ghadimi & Lan, 2013; Duchi et al., 2015; Liu et al., 2020), these methods have recently gained traction for fine-tuning LLMs (Malladi et al., 2023; Deng et al., 2023a; Ling et al., 2024) and have been shown to be able to outperform first-order PEFT methods given enough training time (Zhang et al., 2024b).
However, zeroth-order methods often rely on simple, fixed prompts during fine-tuning. In tasks like sentiment analysis with the SST-2 dataset, templated prompts (e.g., $<CLS>$ text data. It was [terrible $\mid$ great]. $<SEP>$) are crucial for success (Zhang et al., 2024b). These prompts effectively align the text data with task-specific objectives. Therefore, selecting such templates becomes a key hyperparameter, raising the question: Can we automatically discover effective prompts for zeroth-order fine-tuning through prompt tuning? More broadly, can PEFT methods complement zeroth-order fine-tuning for large models? In this work, we propose a new framework to answer this question.
While our focus has thus far been on single-task fine-tuning, many scenarios necessitate multi-task fine-tuning. Multi-task learning (MTL) enables a model to handle multiple tasks simultaneously, fostering knowledge transfer between tasks and improving overall efficiency (Min et al., 2022; Yang et al., 2024). This approach is particularly valuable in low-resource settings, where collecting large labeled datasets can be costly, as is often the case with medical data. In such environments, few-shot learning—where a model is fine-tuned on a high-resource dataset to quickly adapt to new tasks with minimal data—becomes essential (Ye et al., 2021).
To address the challenges of multi-task and few-shot learning in natural language processing, several meta fine-tuning methods have been proposed (Huang et al., 2023; Zhao et al., 2024; Ye et al., 2021; Asadi et al., 2024). However, traditional meta fine-tuning approaches, such as MetaICL (Min et al., 2022), still require full-model first-order gradient calculations, which become computationally expensive with large language models (LLMs) containing billions of parameters. Given the success of zeroth-order methods in fine-tuning LLMs for individual tasks, the potential for adapting their applicability to multi-task learning remains largely unexplored.
With the effectiveness of zeroth-order fine-tuning and the advantages of PEFT for single tasks, a natural question arises: Can we develop a new multi-task and few-shot learning methodology that significantly reduces computational costs while maintaining or even enhancing performance? Specifically, can we leverage the efficiency of zeroth-order fine-tuning alongside the adaptability of PEFT within multi-task and few-shot learning for large language models?
@@ 1.1 Contributions
Bilevel Fine Tuning Framework: In this work we introduce a bilevel framework to enhance zeroth-order fine-tuning for large pre-trained language models using Parameter-Efficient Fine-Tuning (PEFT). It involves two levels of optimization: one for fine-tuning the base model and another for selecting the best PEFT parameters. The two levels are nested such that each level is enhancing the performance of the other level.
Efficient Training Method: We propose the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method to solve the bilevel optimization problem. It incorporates zeroth-order approximations into first-order bilevel methods, avoiding full model gradient computation and addressing the high resource demands of traditional methods, especially for large models. We provide the convergence guarantees of the proposed method. Experiments demonstrate the superiority of Bilevel ZOFO, achieving better results than traditional PEFT and zeroth-order methods in single-task settings while maintaining similar memory efficiency.
Mitigating the sensitivity of hard prompts in ZO fine tuning The standard zeroth-order fine-tuning is very sensitive to the hard prompt used. Our method mitigates this sensitivity problem. In our experiments, the proposed method exhibits significantly lower sensitivity to the choice of hard prompts.
Light-weight Method for Multi-Task Learning: The framework is expanded to multi-task learning, where zeroth-order fine-tuning reduces computational costs, and PEFT models allow efficient task-specific fine-tuning. This results in a lightweight and adaptable meta-training process suitable for new tasks. In the experiments, our method outperforms existing approaches in multi-task learning scenarios.
@@ 2 Related work
2.1 Zeroth order in fine tuning LLMs
MeZO (Malladi et al., 2023) is the first work to use Zeroth-Order (ZO) methods to finetune LLMs for downstream tasks. They demonstrate that their method is compatible with both full-parameter tuning and parameter-efficient tuning techniques, such as LoRA and prefix tuning, while being significantly more computationally efficient. Zhang et al. (2024b) provide a benchmark for ZO optimization in the context of LLM fine-tuning, comparing different ZO optimizers and applying the method to various models. Gautam et al. (2024) introduce variance reduction techniques into ZO methods for fine-tuning, improving both stability and convergence. In addition, ZO methods are applied in federated fine-tuning by Qin et al. (2024) and Ling et al. (2024). Deng et al. (2023b) implement ZO optimization for softmax units in LLMs. Guo et al. (2024b) and Liu et al. (2024b) explore fine-tuning a minimal subset of LLM parameters using ZO methods by sparsifying gradient approximation or the perturbations used in gradient estimation. Tang et al. (2024) investigate the privacy of ZO optimization methods.
2.2 Fine-tuning LLMs for Multitask and Few-Shot Learning
Typical meta-tuning approaches employ First-Order methods to train autoregressive LLMs on a multitask dataset for various tasks (Zhong et al., 2021; Min et al., 2022; Guo et al., 2024a). Zhong et al. (2021) apply meta-training to tasks such as hate speech detection, question categorization, topic classification, and sentiment classification. Guo et al. (2024a) adopt the method from Min et al. (2022) for generating stylistic text. While Min et al. (2022) focus on enhancing the in-context learning ability of the meta-trained model for multitask learning, Zhong et al. (2021) focus on improving zero-shot performance.
During training, Min et al. (2022) sample a task from the dataset for each iteration to perform in-context learning. In contrast to Zhong et al. (2021) and Min et al. (2022), our approach uses a bilevel structure: the full LLM is fine-tuned at the upper level, while parameter-efficient fine-tuning (PEFT) models are tuned at the lower level. At test time, we freeze the meta-tuned base model and fine-tune only the PEFT model using a few-shot setup, which is both more cost-effective and efficient. Crucially, Min et al. (2022) fine tune the full model with first order methods, while we employ a ZO method in meta-tuning the base model at the upper level. Our approach allows us to bypass the need for backpropagation in the meta-model, significantly reducing computational costs.
2.3 Penalty Methods for Bilevel Optimization
Solving a bilevel optimization problem is challenging because the function value in the upper-level objective depends on the optimizer of the lower-level problem. This makes it difficult to compute the gradient of the upper-level objective, also known as the hypergradient. Classical methods require calculating Hessian-vector multiplications to approximate the hypergradient (Franceschi et al., 2017, 2018; Finn et al., 2017; Li et al., 2022; Rajeswaran et al., 2019; Ghadimi & Wang, 2018; Chen et al., 2022; Lorraine et al., 2020). However, when fine-tuning large language models, this process becomes extremely expensive due to the high computational and memory demands.
Recently, new frameworks for bilevel optimization have been introduced (Lu & Mei, 2024; Shen & Chen, 2023; Liu et al., 2024a; Kwon et al., 2023; Liu et al., 2022). These methods bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. The constraint is penalized, allowing the problem to be tackled as a minimax problem using only first-order information. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive.
@@ 2.3 Penalty Methods for Bilevel Optimization
Liu et al. (2024a) and Lu & Mei (2024) explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from Lu & Mei (2024) to approximate part of the upper-level parameters using a ZO approximation In order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
@@
3 Bilevel model and Zeroth-order-first-order method
@@
In this section, we present our bilevel model and the zeroth-order-first-order method for solving it. Let $\mathbf{p}$ represent the parameters of the PEFT model, and $\bm{\theta}$ represent the parameters of the pretrained base model. Given a single downstream task, such as classification, we aim to solve the following optimization problem:
$\min_{{\bm{\theta}}\in\mathbb{R}^{d}}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{f}),$ (1)
where ${\mathbf{p}}\in\mathbb{R}^{d^{\prime}}$ and $F({\bm{\theta}},{\mathbf{p}};\mathcal{B}):=\frac{1}{|\mathcal{B}|}\sum_{x\in\mathcal{B}}F({\bm{\theta}},{\mathbf{p}};x)$ is a loss function given a dataset $\mathcal{B}$.
@@
When ${\mathbf{p}}$ corresponds to the embeddings of the hard prompt (as shown in Table 13 in the appendix of (Malladi et al., 2023)), the model above reduces to classical fine-tuning on a single downstream task. In model (1), the parameters of the PEFT model, ${\mathbf{p}}$, are fixed.
@@
To enhance generalization ability, we split the dataset into two parts: one for tuning the PEFT model (denoted as $\mathfrak{D}_{\mathbf{p}}$) and another for fine-tuning the LLM (denoted as $\mathfrak{D}_{f}$). To maximize performance on downstream tasks, we need the optimal PEFT model parameters that are best suited for the current LLM base model. To achieve this, ${\mathbf{p}}$ should satisfy the following condition:
${\mathbf{p}(\bm{\theta})}\in\mathop{\rm arg\,min}_{{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}}).$
This condition reveals that as the parameters $\theta$ of the LLM change, the parameters ${\mathbf{p}}$ in the PEFT model should also be updated accordingly. Therefore, instead of solving (1), our true objective becomes:
@@
$\begin{split}\min_{{\bm{\theta}}\in\mathbb{R}^{d}}&F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})\\ {\rm{\mathbf{s}}.t.\ }&{\mathbf{p}(\bm{\theta})}\in\mathop{\rm arg\,min}_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}}).\end{split}$ (2)
@@
To solve the bilevel optimization problem (2), classical bilevel methods (as discussed in related work) view (2) as a single-level problem $\min_{\bm{\theta}}F({\bm{\theta}},{\mathbf{p}(\bm{\theta})})$. Since ${\mathbf{p}(\bm{\theta})}$ is the minimizer of another optimization problem, these methods typically require computing the Hessian-vector product (matrix multiplication of $\nabla_{{\bm{\theta}}{\mathbf{p}}}F({\bm{\theta}},{\mathbf{p}})$ and some vector $v$) multiple times to estimate the gradient of $F({\bm{\theta}},{\mathbf{p}(\mathbf{{\bm{\theta}}})})$ with respect to ${\bm{\theta}}$. However, for large language models (LLMs), this approach is computationally prohibitive because the number of parameters in ${\bm{\theta}}$ is too large.
To reduce the computational cost, following Lu & Mei (2024), we consider using a penalty method for the bilevel problem (2). Specifically, (2) is equivalent to the following constrained optimization problem:
$\begin{split}\min_{\bm{\theta}\in\mathbb{R}^{d},\bm{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}&F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{f})\\ {\rm{\mathbf{s}}.t.\ }&F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-\inf_{\mathbf{{\mathbf{s}}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}})\leq 0.\end{split}$ (3)
By penalizing the constraint, we obtain the following penalized problem:
$\min_{\begin{subarray}{c}\bm{\theta}\in\mathbb{R}^{d}\\ \bm{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}\end{subarray}}\!\!F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})+\lambda(F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-\!\!\inf_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}})),$ (4)
where $\lambda>0$. As $\lambda$ increases, the solution to the penalized problem approaches the solution to (3), and thus the solution to (2) (see Lemma 5.4 for an explicit relationship between the stationary points of (4) and those of the original problem (2)). Note that the penalized problem (4) is equivalent to the following minimax problem:
$\min_{{\bm{\theta}}\in\mathbb{R}^{d},{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}\max_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}G_{\lambda}({\bm{\theta}},{\mathbf{p}},{\mathbf{s}}):=\\ F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})+\lambda\big{(}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-F({\bm{\theta}},{\mathbf{s}};\mathfrak{D}_{\mathbf{p}})\big{)}.$ (5)
In this way, we can solve the bilevel problem as a minimax problem. The basic minimax algorithm works as follows: at iteration $k$, we first solve the maximization problem $\max_{\mathbf{{\mathbf{s}}}}G_{\lambda}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}})$ with $({\bm{\theta}}^{k},{\mathbf{p}}^{k})$ fixed. For example, we can update ${\mathbf{s}}^{k}$ using an inner loop with stochastic gradient descent (SGD). Let ${\mathbf{s}}^{k+1}$ be the result of this inner loop. Then, in the outer loop, we update $({\bm{\theta}}^{k},{\mathbf{p}}^{k})$ by solving $\min_{{\bm{\theta}},{\mathbf{p}}}G_{\lambda}({\bm{\theta}},{\mathbf{p}},{\mathbf{s}}^{k+1})$ with ${\mathbf{s}}^{k+1}$ fixed. Again, SGD can be used to update ${\bm{\theta}}^{k}$ and ${\mathbf{p}}^{k}$. The conceptual algorithm is presented in Algorithm 1.

### v2

@@ Abstract
Fine-tuning pre-trained Large Language Models (LLMs) for downstream tasks using First-Order (FO) optimizers presents significant computational challenges. Parameter-Efficient Fine-Tuning (PEFT) methods address these by freezing most model parameters and training only a small subset. However, PEFT often underperforms compared to full fine-tuning when high task-specific accuracy is required. Zeroth-Order (ZO) methods fine-tune the entire pre-trained model without back-propagation, estimating gradients through forward passes only. While memory-efficient, ZO methods suffer from slow convergence and high sensitivity to prompt selection. We bridge these two worlds with Bilevel-ZOFO, a bilevel optimization method that couples fast, local FO-PEFT adaptation at the inner level with stable, memory-efficient ZO updates of the full backbone at the outer level. The FO-PEFT inner loop performs fast, low-memory local adaptation that reduces the variance of ZO estimates and stabilizes the search, guiding the outer ZO updates of the full backbone and reducing prompt sensitivity. In the mean time, the outer ZO provides better generalization ability for PEFT. We provide theoretical convergence guarantees and empirically demonstrate that Bilevel-ZOFO significantly outperforms existing ZO and FO-PEFT methods, achieving 2–4× faster training while maintaining similar memory efficiency. Additionally, we show by updating the backbone with ZO and adapting only a tiny FO-PEFT block per task, Bilevel-ZOFO combines full-model capacity with few-shot efficiency, making it a very efficient meta-learning algorithm that quickly adapts to new tasks.
@@ 1 Introduction
Fine-tuning pretrained Large Language Models (LLMs) has become a standard approach for downstream tasks. Traditional First-Order (FO) optimizers like Adam [20], commonly used for this process, rely on backpropagation. However, as highlighted in Malladi et al. [33], computing gradients for LLMs can require up to 12 times the memory needed for inference. This scaling challenge becomes even more pronounced as models grow larger, imposing significant memory demands and complicating the fine-tuning process, especially in resource-constrained environments.
@@
To address these computational challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have been developed. These techniques freeze most of the model’s parameters and train only a small subset, significantly reducing both memory and computational overhead. Popular PEFT approaches include prompt tuning, LoRA fine-tuning, and prefix tuning. Prompt tuning [22, 38, 56, 46] optimizes continuous prompt vectors that are concatenated with the input embeddings, while prefix tuning [24] introduces learnable prefix tokens that serve as conditioning variables at each transformer layer. LoRA (Low-Rank Adaptation) [16, 15] modifies the model’s attention and feedforward layers by injecting low-rank trainable matrices, further reducing the resources required for fine-tuning.
[…]
@@
While Parameter-Efficient Fine-Tuning (PEFT) methods reduce training costs and memory usage, they may not always achieve the same level of task-specific performance as full model fine-tuning. Research has shown that for tasks requiring high accuracy, complex adaptations, or domain-specific knowledge, full fine-tuning often outperforms PEFT approaches due to its ability to adjust all model parameters for better adaptation [16, 24, 57]. To make full model fine-tuning more computationally feasible, Zeroth-Order (ZO) methods offer an alternative by reducing the high computational cost. Rather than computing gradients via backpropagation, zeroth-order methods estimate the gradient using only the forward pass. Initially explored in the 1990s [48, 36, 11, 6, 28], these methods have recently gained traction for fine-tuning LLMs [33, 4, 25] and have been shown to be able to outperform FO PEFT methods given enough training time [61].
A major limitation of ZO methods is their slow convergence due to the need for gradient estimation. For instance, MeZO [33] required 10 times more iterations than PEFT baselines to match or exceed their performance. Additionally, ZO methods suffer from extreme sensitivity to prompt selection. In tasks like sentiment analysis with the SST-2 dataset, templated prompts (e.g., <CLS><CLS> text data. It was [terrible $\mid$ great]. <SEP><SEP>") are crucial for success [61]. These prompts effectively align the text data with task-specific objectives. As a result, prompt selection becomes an important hyperparameter that can significantly affect performance. In particular, ZO methods have been shown to be highly sensitive to prompt selections  [33]. Without proper prompts, the performance of MeZO can drastically drop (Table 1).
In this paper, we ask: Can zeroth-order (ZO) and PEFT methods be smoothly integrated to mutually enhance each other—achieving greater robustness to prompt variations, faster convergence, and better performance than either method alone—while maintaining memory efficiency comparable to each individually? We target settings where (i) full FO fine-tuning is impractical due to memory/throughput, (ii) pure PEFT lacks full-model capacity on harder adaptations, and (iii) pure ZO is slow and highly prompt-sensitive.
We propose Bilevel-ZOFO, a novel bilevel optimization framework explicitly designed to leverage the complementary strengths of these two approaches:
  • At the inner level, FO-PEFT rapidly performs targeted, local adaptation using first-order gradients, stabilizing training and mitigating sensitivity to task-specific prompts that ZO methods need.
  • At the outer level, a ZO method updates the full backbone model parameters efficiently, guided by the stable and informative inner-level adaptation. This full model finetuning enhances the model’s generalization ability, enables a more sophisticated understanding, and improves transfer to new tasks.
This clear separation enables efficient bilevel optimization, addressing the major drawbacks of pure ZO methods (slow convergence, prompt sensitivity) and pure PEFT methods (limited full-model adaptation). Extensive ablation studies empirically verify this synergy, demonstrating faster convergence and more robust performance.
@@ 1.1 Contributions
We summarize our main contributions as follows:
  1. 1.
    We propose Bilevel-ZOFO, a theoretically grounded and practical bilevel optimization method that enhances zeroth-order (ZO) optimization by selecting the best prompt and thereby improves ZO fine-tuning with first-order PEFT (FO-PEFT). At the same time, it strengthens PEFT by leveraging full fine-tuning through ZO updates.
  2. 2.
    Bilevel-ZOFO reduces ZO sensitivity to prompt choices and significantly accelerates convergence, achieving state-of-the-art performance with minimal memory overhead.
  3. 3.
    Extensive experiments confirm that Bilevel-ZOFO consistently outperforms existing FO-PEFT and ZO baselines across diverse tasks.
  4. 4.
    By updating the backbone with ZO and adapting only a tiny FO-PEFT block per task, our method couples full-capacity transfer with few-shot efficiency. We show that this design has strong potential for efficient meta-learning, demonstrating improved multi-task adaptation with minimal computational resources.
@@ 2 Related work
2.1 Zeroth-Order Methods in Fine-Tuning LLMs
MeZO [33] pioneered zeroth-order (ZO) fine-tuning for LLMs, demonstrating compatibility with both full-model and PEFT approaches while improving computational efficiency. Subsequent work provided a benchmark for ZO optimization methods [61], and expanded ZO applications to variance reduction [10], federated fine-tuning [39, 25], softmax layers [5], sparse tuning [14, 29], and privacy [49]. In contrast, we propose a bilevel training framework that unifies the strengths of ZO full-model and FO PEFT fine-tuning, outperforming both individually, while using as much resources.
2.2 First-Order Methods for Bilevel Optimization
@@
Bilevel optimization is computationally demanding, especially for LLMs, due to the cost of computing hypergradients. Classical approaches rely on second-order methods [8, 9, 7, 23, 41, 12, 2, 30], while recent work [31, 44, 27, 21, 26, 18, 32] bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. We build on this by incorporating ZO approximations into the upper-level optimization to bypass full gradient computation for LLMs. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive. Liu et al. [27] and Lu and Mei [31] explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from Lu and Mei [31] to approximate part of the upper-level parameters using a ZO approximation in order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
3 Bilevel Model with Zeroth-Order-First-Order Method
@@
In this section, we introduce our bilevel model and the zeroth-order-first-order method for solving it.
3.1 Preliminaries and Notation
@@
Let $\mathbf{p}\in\mathbb{R}^{d^{\prime}}$ represent the parameters of the PEFT model, and $\bm{\theta}\in\mathbb{R}^{d}$ represent the parameters of the pretrained base model. We denote the loss function given a dataset $\mathcal{D}$ as $F({\bm{\theta}},{\mathbf{p}};\mathcal{D}):=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}F({\bm{\theta}},{\mathbf{p}};x)$. Given a single downstream task, such as classification, we aim to solve the following optimization problem:
$\min_{{\bm{\theta}}\in\mathbb{R}^{d}}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}).$ (1)
@@
Where $\mathbf{p}$} corresponds to the embeddings of the hard prompt (as shown in Table 13 in the appendix of [33]), the model above reduces to classical fine-tuning on a single downstream task. In model (1), the parameters of the PEFT model, $\mathbf{p}$}, are fixed.
@@
To enhance generalization ability, we split the dataset $\mathfrak{D}$ into two parts: one for tuning the PEFT model (denoted by $\mathfrak{D}_{\mathbf{p}}$) and another for fine-tuning the LLM (denoted by $\mathfrak{D}_{f}$). To maximize performance on downstream tasks, we need the optimal PEFT model parameters that are best suited for the current LLM base model. To achieve this, we require $\mathbf{p}$} to satisfy the following condition:
${\mathbf{p}}\in\mathop{\rm arg\,min}_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{s}};\mathfrak{D}_{\mathbf{p}}).$
[…]
@@
$\begin{split}\min_{{\bm{\theta}}\in\mathbb{R}^{d}}F({\bm{\theta}},\mathbf{p};\mathfrak{D}_{f})\ {\rm s.\ t.}\ {\mathbf{p}}\in\mathop{\rm arg\,min}_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}}).\end{split}$ (2)
@@
3.2 Bilevel Model
Eq. (2) is an instance of a bilevel optimization problem. To solve it, classical bilevel methods (as discussed in related work) view Eq. (2) as a single-level problem $\min_{\bm{\theta}}F({\bm{\theta}},{\mathbf{p}})$. Since $\mathbf{p}$} is the minimizer of another optimization problem, these methods typically require computing the Hessian-vector product (matrix multiplication of $\nabla_{{\bm{\theta}}{\mathbf{p}}}F({\bm{\theta}},{\mathbf{p}})$ and some vector $v$) multiple times to estimate the gradient of $F({\bm{\theta}},\mathbf{p})$ with respect to $\bm{\theta}$}. However, for large language models (LLMs), this approach is computationally prohibitive because the number of parameters in $\bm{\theta}$} is too large.
To reduce the computational cost, following [31], we consider using a penalty method for the bilevel problem (2). Specifically, (2) is equivalent to the following constrained optimization problem:
$\begin{split}\min_{\bm{\theta}\in\mathbb{R}^{d},\bm{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{f})\ {\rm{\mathbf{s}}.t.\ }\ F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-\inf_{\mathbf{{\mathbf{s}}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}})\leq 0.\end{split}$ (3)
By penalizing the constraint with a constant $\lambda>0$, we obtain the following penalized problem:
$\min_{\begin{subarray}{c}\bm{\theta}\in\mathbb{R}^{d}\\ \bm{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}\end{subarray}}\!\!F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})+\lambda(F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-\!\!\inf_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}})).$ (4)
As $\lambda$ increases, the solution to the penalized problem approaches the solution to (3), and thus the solution to (2) (see Lemma B.4 for an explicit relationship between the stationary points of (4) and those of the original problem (2)). Note that the penalized problem (4) is equivalent to the following minimax problem:
$\min_{{\bm{\theta}}\in\mathbb{R}^{d},{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}\max_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}G_{\lambda}({\bm{\theta}},{\mathbf{p}},{\mathbf{s}}):=F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})+\lambda\big(F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-F({\bm{\theta}},{\mathbf{s}};\mathfrak{D}_{\mathbf{p}})\big).$ (5)